# CEFR-Annotated WordNet: LLM-Based Proficiency-Guided Semantic Database for Language Learning

### v1

@@ 5.1. Datasets and Experimental Settings
To train our classifiers, we needed a corpus with CEFR-level annotations for words in context. As the original SemCor corpus lacks this information, we created the “SemCor-CEFR corpus” by assigning CEFR levels to its senses using our annotated WordNet. Table 2 summarizes the word distributions in the EVP Online examples (dictionary and learner examples combined) and our SemCor-CEFR corpus. While our corpus has fewer word types (#types) than the EVP examples, it contains substantially more word instances (#words) and reflects a more natural, imbalanced distribution of proficiency levels. For our experiments, we used 10% of the EVP examples as the test set, whereas the remaining 90% and the SemCor-CEFR corpus were used for training and validation. The task for each classifier was to predict the CEFR level (a six-way classification) of a target word within a given sentence. We report the F1 score for each level, along with Macro-F1 and Micro-F1 scores for overall performance.
@@ 5.2. Classifiers
We compared the performance of a baseline model, ME6 Contextual, with several LLM-based approaches: zero-shot, few-shot, and fine-tuned models.
@@
We reimplemented ME6 Contextual as a baseline. This method uses BERT-based contextual embeddings to train a support vector classifier (SVC) that predicts CEFR levels. For the hyperparameters of BERT and SVC that were not explicitly specified in the original study Aleksandrova and Pouliot (2023), we used their default settings. Although this model supports MWEs, we exclude MWEs to align with the scope of WordNet, which contains almost none. We trained three versions of this model: one on 90% of the EVP examples, one on our SemCor-CEFR corpus, and one on a mixture of both. As noted in Section 4, a single sense in our corpus may have multiple levels; therefore, we created one training example per level when training on our data. If the model trained on our corpus performs comparably to the one trained on the gold-standard EVP examples, it would support the accuracy of our level annotations.
@@
We evaluated an LLM’s inherent ability to classify CEFR levels without any examples. Using the prompt and parameter settings in Appendix 13.1, we provided the model (GPT-5, checkpoint gpt-5-2025-08-07) with a target word and its context and asked it to output the corresponding CEFR level.
Few-Shot Trained LLMs.
We also evaluated the LLM’s performance with 6- and 18-shot prompting, using the template provided in Appendix 13.1. This prompt provides the model with training examples to serve as clues for classifying a word sense in its context. For the 6-shot setting, we provided one training example for each of the six CEFR levels (i.e., examples), whereas for the 18-shot setting, we used three examples per level (i.e., examples). The target words and their usage examples were randomly selected from the 90% of EVP examples reserved for training. The LLM and parameter settings were the same as those used for the zero-shot experiments.
@@
We fine-tuned the lightweight and cost-effective GPT-4.1 mini model (checkpoint gpt-4.1-mini-2025-04-14) on three different datasets: 90% of the EVP examples, our SemCor-CEFR corpus, and a mixture of both. Similar to the ME6 Contextual baseline evaluation, the high accuracy of our annotations would be confirmed if the model fine-tuned on our corpus achieved comparable or better performance than that trained on the gold-standard EVP examples. We also trained a model on the mixed corpus to test for any synergistic effects. Senses with multiple CEFR levels in our corpus were treated as separate training examples for each level. For fine-tuning, we used a 90%/10% split for training and validation. The training data were formatted by populating the zero-shot template from Appendix 13.1 with each target word and sentence, using the corresponding CEFR level as the correct answer. The default (auto) hyperparameters used for fine-tuning are listed in Appendix 13.1.
@@
For words whose CEFR level is unambiguous in the EVP Online (i.e., all senses share the same level), running a full six-level classification is computationally inefficient and increases the risk of classification errors. To address this, we developed a hybrid approach. We first built a knowledge base—a word-level list from the EVP Online containing only words with a single CEFR level. For each target word, we first checked this list. If the word was present, we directly assigned its recorded level; otherwise, we used one of the fine-tuned LLMs for classification. We applied this method to the LLMs fine-tuned on the EVP examples, our SemCor-CEFR corpus, and the mixed corpus to observe the differences in classification accuracy.
@@ 5.3. Results
Table 3 reports the F1 scores for each classifier. In this table, FT denotes the fine-tuned LLMs, and FT+KB refers to the fine-tuned LLMs combined with the knowledge-based approach. The training datasets used are EVP (90% of the EVP examples), SemCor-CEFR (our annotated SemCor corpus), and Mixture (a combination of both). Because the class distribution in our data was imbalanced, we used the Macro-F1 score as the primary metric for evaluating overall performance, as it assigns equal weight to each class and thus mitigates the effects of frequency imbalance.
@@
The ME6 Contextual classifier achieved a Macro-F1 score of at least 0.5 across all training sets. However, its performance on the SemCor-CEFR corpus was 0.08 points lower than that on the EVP data. We attribute this gap to the model’s vector construction method, which averages the vectors for all instances of a given word and CEFR level, resulting in a single vector per word-level pair. As shown in Table 2, our SemCor-CEFR corpus has fewer unique word types than the EVP data. Consequently, despite having a higher total word frequency, it yields fewer training vectors, which likely contributed to the performance drop. Consistent with this interpretation, the classifier trained on the Mixture dataset, which contained the most training examples, achieved the best performance among the ME6 Contextual models.
@@
The zero-shot LLM achieved a Macro-F1 of 0.42, the lowest score among all methods and well below that of the ME6 Contextual baseline. Its F1 scores for the C1 and C2 levels were particularly low (below 0.3), indicating that the LLM’s internal knowledge alone is insufficient for classifying advanced-level senses. Providing in-context examples via few-shot prompting raised the Macro-F1 score to 0.47 (6-shot) and 0.48 (18-shot). This improvement, which aligns with prior findings Enomoto et al. (2024); Smǎdu et al. (2024), stemmed from supplementing the model’s knowledge of C1 and C2 senses. Nevertheless, the performance of the few-shot models remained significantly lower than that of ME6 Contextual.
@@
Fine-tuning proved to be a highly effective approach for developing LLM classifiers, improving the Macro-F1 score by at least 0.17 points over the few-shot methods. Notably, the FT model trained on our SemCor-CEFR corpus performed comparably to that trained on the gold-standard EVP data, despite being optimized on an entirely different dataset from the test set. Moreover, an analysis of its errors (Figure 6 in Appendix 13.2) shows that misclassifications, especially for C1-level senses, were often assigned to adjacent proficiency levels, which would minimize confusion for learners. This strong performance is likely because the LLM could be trained directly on the rich and varied usage examples in the SemCor-CEFR corpus. The model trained on the Mixture dataset achieved a Macro-F1 of 0.73. These results confirm both the high accuracy of the CEFR annotations in our WordNet and the effectiveness of combining the EVP and SemCor-CEFR corpora.
@@
The hybrid FT+KB approach, which combines fine-tuned LLMs with a knowledge base, yielded the best performance. This method improved the Macro-F1 score by 0.08–0.13 points over the FT models alone, with a consistent performance trend across the different training sets. The classifier trained on the Mixture dataset achieved the highest F1 scores across all levels, exceeding 0.8 for every level except B1 and C1. This suggests that a substantial portion of the test set comprises words with unambiguous CEFR levels. In such cases, the knowledge base can handle the classification without LLM inference, improving both accuracy and computational efficiency.
@@ 6. Discussion
While our FT and FT+KB classifiers trained on EVP examples demonstrated strong performance, these results may be inflated, as both the fine-tuning and test sets were drawn from the same source. For real-world applications, a CEFR-level classifier must be effective across various domains, not just dictionary and learner examples. However, gold-standard, sense-level CEFR annotations for diverse corpora are scarce. To address this, we evaluated our classifiers’ generalizability by examining the correlation between their predicted CEFR levels and the lexical complexity scores in the CompLex 2.0 dataset. This dataset, used for the LCP task, spans three distinct genres—Europarl, the Bible, and biomedical texts— and contains target words rated by multiple annotators on a continuous complexity scale from 0 to 1. We used our classifiers to predict CEFR levels (mapped as integers 1–6) for 7,662 target words in the CompLex 2.0 training set and then computed the Spearman rank correlation coefficient between our predictions and the dataset’s complexity scores. We did not expect a very high correlation, as complexity scores are continuous while CEFR levels are discrete.
@@
A notable finding from the results presented in Table 4 is that classifiers trained on the EVP examples, despite their high accuracy on the EVP test set, showed significantly low correlation with the CompLex 2.0 scores. This suggests that models fine-tuned solely on EVP data may have learned superficial, dataset-specific cues and therefore exhibit poor generalizability to other text genres. By contrast, classifiers trained on our SemCor-CEFR corpus achieved correlation coefficients exceeding 0.5, indicating a moderate relationship between their predictions and lexical complexity. We attribute this significant improvement to the broad genre coverage of the SemCor corpus combined with the high accuracy of our CEFR-level annotations. Consequently, classifiers fine-tuned on our corpus are better suited for use with educational materials from diverse sources.
@@ 6.2. Implications for L2 Learners
Our findings have important implications for L2 learners, who often struggle with the fine-grained sense distinctions in WordNet. By annotating WordNet senses with CEFR levels and integrating them into resources such as the SemCor-CEFR corpus, our method helps align lexical information with learner proficiency and pedagogical needs. Although the practical benefits of this approach require empirical validation through classroom or longitudinal studies, it offers two key advantages. First, it allows learners to focus on senses that match their proficiency level, reducing the cognitive load associated with more advanced or nuanced meanings. Second, our high-performing classifier (Macro-F1 of 0.81) can be integrated into educational tools to quickly identify complex lexical items in a text, enabling immediate scaffolding. The model’s superior performance on levels A1 through B2 is particularly beneficial for beginner and intermediate learners who are building their foundational vocabulary. These advancements can allow educators and autonomous learners to adopt more adaptive and efficient strategies for vocabulary instruction. However, further research is required to confirm whether these benefits persist across diverse learning environments and over extended periods.
@@ 7. Conclusions
In this study, we introduced an LLM-based method for annotating WordNet senses with CEFR levels and used it to construct a CEFR-annotated WordNet. This new resource provides 10,995 proficiency annotations for 10,644 senses across 5,645 lemmas. Using this annotated WordNet, we also created the SemCor-CEFR corpus, a large-scale resource containing over 110,000 sense-level CEFR annotations. To validate our approach, we trained contextual lexical CEFR-level classifiers on our corpus and found that they performed comparably to those trained on gold-standard data. Moreover, by combining our corpus with the gold-standard levels, we developed a practical classifier that achieves a Macro-F1 score of 0.81, confirming the high accuracy and utility of our CEFR annotations. Our analysis also showed that the predictions from our classifiers correlate with the lexical complexity scores in the CompLex 2.0 dataset, suggesting that they can generalize effectively across diverse text genres.
This work is part of the “Learner’s WordNet Project,” which aims to integrate NLP methods—particularly WordNet’s rich semantic network—with educational technology to support more efficient and effective L2 learning. Future work will focus on expanding the CEFR-level coverage in our WordNet, evaluating its pedagogical effectiveness in real-world learning scenarios, and developing related applications. To support this expansion, we plan to build a classifier that can accurately assign CEFR levels to previously unannotated word senses.
@@ 8. Acknowledgements
This work was supported in part by JSPS KAKENHI Grant Numbers JP22K02825, JP22K18006, JP25K21351, and JP24K03052. This publication/presentation/research report has made use of the English Vocabulary Profile. This resource is based on extensive research using the Cambridge Learner Corpus and is part of the English Profile programme, which aims to provide evidence about language use that helps to produce better language teaching materials. See https://englishprofile.org/ for more information.
@@ 9. Ethical Considerations
We accessed the EVP Online strictly under the EVP website’s Terms of Use444https://englishprofile.org/?menu=evp-terms-of-use and Cambridge University Press’s text and data mining (TDM) terms555https://www.cambridge.org/us/legal/copyright. Any EVP content temporarily cached to local storage for this project was deleted upon the project’s completion. All released artifacts are built exclusively from WordNet and SemCor, whose licenses permit copying, modification, and redistribution. No EVP data (including entries, examples, glosses, or metadata) are included in any of the released resources.
For verification, we employed proprietary LLMs from OpenAI and enabled the opt‑out setting to ensure that data were not used for model training. However, in any downstream use cases that involve personal or sensitive data, we recommend deploying open source LLMs in a local environment to reduce the risk of unintended disclosure. The resources introduced in this work are compatible with such locally hosted LLMs. The CEFR levels provided by our resources are model-based estimates and must not be used as the sole criterion for high-stakes educational decisions, such as promotion, pass/fail judgments, or selective admissions.

### v2

@@ 5.1. Datasets and Experimental Settings
To train our classifiers, we required a corpus with CEFR-level annotations for words in context. Because the original SemCor corpus does not include this information, we created the “SemCor-CEFR corpus” by assigning CEFR levels to its senses using our annotated WordNet. Table 2 summarizes the word distributions in the EVP Online examples, combining dictionary and learner examples, and in the SemCor-CEFR corpus. Although our corpus contains fewer word types (# types) than the EVP examples, it includes substantially more word instances (# words) and reflects a more natural and imbalanced distribution of proficiency levels. For our experiments, we used 10% of the EVP examples as the test set. For training and validation, depending on the setting, we used (i) the remaining 90% of the EVP data, (ii) the SemCor-CEFR corpus, or (iii) a combination of both. The task for each classifier was to predict the CEFR level, formulated as a six-way classification problem, of a target word within a given sentence. We report the F1 score for each level, together with Macro-F1 and Micro-F1 scores to measure overall performance.
@@ 5.2. Classifiers
We compared the performance of a baseline model, ME6 Contextual, with several LLM-based approaches, including zero-shot, few-shot, and fine-tuned models.
@@
We reimplemented ME6 Contextual as a baseline. This method uses BERT-based contextual embeddings to train a support vector classifier (SVC) that predicts CEFR levels. For the hyperparameters of BERT and SVC that were not explicitly specified in the original study Aleksandrova and Pouliot (2023), we used their default settings. Although the model supports MWEs, we excluded them to align with the scope of WordNet, which contains very few MWEs. We trained three versions of the model: one on 90% of the EVP examples, one on the SemCor-CEFR corpus, and one on a combination of both. If the model trained on our corpus performs comparably to the model trained on the gold-standard EVP examples, this would support the accuracy of our level annotations.
As noted in Section 4, a single sense in our corpus may be associated with multiple CEFR levels. Therefore, when training on our data, we created one training example per level. These multi-level assignments arise when a single WordNet sense aligns with multiple, more fine-grained EVP senses. Rather than discarding them as noise, we split them into separate training instances—one per level—and treat them as alternative supervision signals. Because 96.84% of senses in our WordNet receive a single CEFR label (Table 6 in Appendix 13.2), we expect the overall impact of this multi-label handling to be limited. Nevertheless, since the EVP evaluation examples provide a single gold level per sentence, this strategy may introduce label ambiguity and reduce performance. Importantly, it should not provide an advantage over models trained solely on EVP data.
@@
We evaluated the LLM’s inherent ability to classify CEFR levels without providing any examples. Using the prompt and parameter settings described in Appendix 13.1, we provided the model (GPT-5, checkpoint gpt-5-2025-08-07) with a target word and its context and asked it to output the corresponding CEFR level.
Few-Shot Prompted LLMs.
We also evaluated the LLM’s performance under 6-shot and 18-shot prompting, using the template provided in Appendix 13.1. This prompt provides the model with training examples to serve as clues for classifying a word sense in its context. In the 6-shot setting, we provided one training example for each of the six CEFR levels, that is, examples. In the 18-shot setting, we used three examples per level, that is, examples. The target words and usage examples were randomly selected from the 90% of EVP examples reserved for training. The LLM and parameter settings were identical to those used in the zero-shot experiments.
@@
We fine-tuned the lightweight and cost-effective GPT-4.1 mini model (checkpoint gpt-4.1-mini-2025-04-14) on three datasets: 90% of the EVP examples, the SemCor-CEFR corpus, and a combination of both. As in the evaluation of the ME6 Contextual baseline, the accuracy of our annotations would be supported if the model fine-tuned on our corpus achieved performance comparable to or better than that of the model trained on the gold-standard EVP examples. We also trained a model on the combined corpus to examine potential synergistic effects. Senses associated with multiple CEFR levels in our corpus were treated as separate training examples for each level. For fine-tuning, we used a 90%/10% split for training and validation. The training data were formatted by filling the zero-shot template in Appendix 13.1 with each target word and sentence, using the corresponding CEFR level as the correct label. The default auto hyperparameters used for fine-tuning are listed in Appendix 13.1.
@@
For words whose CEFR level is unambiguous in the EVP Online, that is, all senses share the same level, performing a full six-level classification is computationally inefficient and increases the risk of errors. To address this issue, we developed a hybrid approach. We first constructed a knowledge base—a word-level list derived from the EVP Online that includes only words associated with a single CEFR level. For each target word, we checked this list first. If the word was present, we directly assigned its recorded level. Otherwise, we used one of the fine-tuned LLMs for classification. We applied this method to the LLMs fine-tuned on the EVP examples, the SemCor-CEFR corpus, and the combined corpus to compare differences in classification accuracy.
@@ 5.3. Results
Table 3 reports the F1 scores for each classifier. In the table, FT denotes the fine-tuned LLMs, and FT+KB refers to the fine-tuned LLMs combined with the knowledge-based approach. The training datasets used are EVP (90% of the EVP examples), SemCor-CEFR (our annotated SemCor corpus), and Mixture (a combination of both). Because the class distribution in our data is imbalanced, we use the Macro-F1 score as the primary metric for overall evaluation, as it assigns equal weight to each class and mitigates the effects of frequency imbalance.
@@
The ME6 Contextual classifier achieved a Macro-F1 score of at least 0.5 across all training sets. However, its performance on the SemCor-CEFR corpus was 0.08 points lower than that on the EVP data. We attribute this gap to the model’s vector construction method, which averages the vectors for all instances of a given word and CEFR level to produce a single vector for each word-level pair. As shown in Table 2, our SemCor-CEFR corpus has fewer unique word types than the EVP data. Consequently, despite having a higher total word frequency, it yields fewer training vectors, which likely contributed to the performance drop. Consistent with this interpretation, the classifier trained on the Mixture dataset, which included the largest number of training examples, achieved the best performance among the ME6 Contextual models.
@@
The zero-shot LLM achieved a Macro-F1 score of 0.42, the lowest among all methods and well below that of the ME6 Contextual baseline. Its F1 scores for the C1 and C2 levels were particularly low, below 0.3, indicating that the LLM’s internal knowledge alone is insufficient for classifying advanced-level senses. Providing in-context examples through few-shot prompting increased the Macro-F1 score to 0.47 in the 6-shot setting and 0.48 in the 18-shot setting. This improvement, consistent with prior findings Enomoto et al. (2024); Smǎdu et al. (2024), resulted from supplementing the model’s knowledge of C1 and C2 senses. Nevertheless, the performance of the few-shot models remained substantially lower than that of ME6 Contextual.
@@
Fine-tuning proved to be a highly effective approach for developing LLM classifiers, improving the Macro-F1 score by at least 0.17 points compared with the few-shot methods. Notably, the FT model trained on the SemCor-CEFR corpus performed comparably to the model trained on the gold-standard EVP data, despite being optimized on a dataset entirely different from the test set. Moreover, an analysis of its errors (Figure 6 in Appendix 13.3) shows that misclassifications, particularly for C1-level senses, were often assigned to adjacent proficiency levels, which would likely minimize confusion for learners. This strong performance is likely due to the model being fine-tuned on the rich and varied usage examples in the SemCor-CEFR corpus. The model trained on the Mixture dataset achieved a Macro-F1 score of 0.73. These results provide indirect evidence supporting the quality of the CEFR annotations in our WordNet and effectiveness of combining the EVP and SemCor-CEFR corpora.
@@
The hybrid FT+KB approach, which combines fine-tuned LLMs with a knowledge base, achieved the best overall performance. This method improved the Macro-F1 score by 0.08 to 0.13 points compared with the FT models alone, with a consistent trend across training sets. The classifier trained on the Mixture dataset achieved the highest F1 scores at all levels, exceeding 0.8 for every level except B1 and C1. This pattern suggests that a substantial portion of the test set consists of words with unambiguous CEFR levels. In such cases, the knowledge base can assign the correct level without relying on LLM inference, thereby improving both accuracy and computational efficiency.
@@ 6. Discussion
Although the FT and FT+KB classifiers trained on EVP examples demonstrated strong performance, these results may be inflated because both the fine-tuning and test sets were drawn from the same source. For real-world applications, a CEFR-level classifier must perform well across diverse domains, not only on dictionary and learner examples. However, gold-standard sense-level CEFR annotations for heterogeneous corpora are scarce. To examine generalizability, we used an indirect proxy by analyzing the correlation between the predicted CEFR levels and lexical complexity scores in the CompLex 2.0 dataset. This dataset, developed for the LCP task, spans three genres—Europarl, the Bible, and biomedical texts—and contains target words rated by multiple annotators on a continuous complexity scale from 0 to 1. We applied our classifiers to predict CEFR levels, mapped to integers 1 to 6, for 7,662 target words in the CompLex 2.0 training set. We then computed the Spearman rank correlation coefficient between the predicted levels and the dataset’s complexity scores. We did not expect a high correlation because complexity scores are continuous, whereas CEFR levels are discrete.
@@
A notable finding in Table 4 is that classifiers trained on the EVP examples, despite achieving high accuracy on the EVP test set, exhibited very low correlation with the CompLex 2.0 scores. This suggests that models fine-tuned solely on EVP data may have learned superficial, dataset-specific cues, with limited transfer to other genres. By contrast, classifiers trained on the SemCor-CEFR corpus achieved correlation coefficients above 0.5, indicating a moderate relationship between their predictions and lexical complexity. We attribute this improvement to the broad genre coverage of the SemCor corpus together with the high quality of our CEFR-level annotations. These findings indicate that classifiers fine-tuned on our corpus are better suited for application to educational materials drawn from diverse sources.
@@ 6.2. Implications for L2 Learners
Our findings have important implications for L2 learners, who often struggle with the fine-grained sense distinctions in WordNet. By annotating WordNet senses with CEFR levels and integrating them into resources such as the SemCor-CEFR corpus, our approach aligns lexical information more closely with learner proficiency and pedagogical needs. Although the practical benefits of this approach require empirical validation through classroom-based or longitudinal studies, it offers two main advantages. First, it allows learners to focus on senses that match their proficiency level, reducing the cognitive load associated with more advanced or nuanced meanings. Second, our high-performing classifier (Macro-F1 of 0.81) can be integrated into educational tools to quickly identify complex lexical items in a text, enabling immediate scaffolding. The model’s strong performance on levels A1 through B2 is particularly beneficial for beginner and intermediate learners who are building foundational vocabulary. These advances may enable educators and independent learners to adopt more adaptive and efficient strategies for vocabulary instruction. However, further research is needed to determine whether these benefits persist across diverse learning environments and over extended periods.
@@ 7. Conclusions
In this study, we introduced an LLM-based method for annotating WordNet senses with CEFR levels and used it to construct a CEFR-annotated WordNet. This resource provides 10,995 proficiency annotations for 10,644 senses across 5,645 lemmas. Using the annotated WordNet, we also created the SemCor-CEFR corpus, a large-scale resource containing more than 110,000 sense-level CEFR annotations. To validate our approach, we trained contextual lexical CEFR-level classifiers on this corpus and found that they performed comparably to models trained on gold-standard data. Furthermore, by combining our corpus with gold-standard levels, we developed a practical classifier that achieves a Macro-F1 score of 0.81, providing indirect evidence of the utility and consistency of our CEFR annotations. Our analysis also showed that the predictions generated by our classifiers correlate with the lexical complexity scores in the CompLex 2.0 dataset, suggesting moderate alignment with lexical complexity across diverse text genres.
This work is part of the “Learner’s WordNet Project,” which seeks to integrate NLP methods—particularly WordNet’s rich semantic network—with educational technology to support more efficient and effective L2 learning. Future work will focus on expanding CEFR-level coverage in our WordNet, evaluating its pedagogical effectiveness in real-world learning contexts and developing related applications. To support this effort, we plan to build a classifier capable of accurately assigning CEFR levels to previously unannotated word senses.
@@ 8. Acknowledgements
This work was supported in part by JSPS KAKENHI Grant Numbers JP22K02825, JP22K18006, JP25K21351, and JP24K03052. This publication/presentation/research report makes use of the English Vocabulary Profile. This resource is based on extensive research using the Cambridge Learner Corpus and forms part of the English Profile program, which aims to provide evidence about language use to support the development of improved language teaching materials. See https://englishprofile.org/ for more information.
@@ 9. Ethical Considerations
We accessed the EVP Online strictly in accordance with the EVP website’s Terms of Use555https://englishprofile.org/?menu=evp-terms-of-use and Cambridge University Press’s text and data mining (TDM) terms666https://www.cambridge.org/us/legal/copyright. Any EVP content temporarily cached to local storage for this project was deleted upon the project’s completion. All released artifacts are built on WordNet and SemCor, whose licenses permit copying, modification, and redistribution, and include only derived CEFR labels mapped to WordNet sense keys. No verbatim EVP content, including entries, examples, glosses, or metadata, is included in the released resources. The EVP-derived word-level lookup list used in the FT+KB analysis (Section 5.2) was used solely for internal evaluation and is not redistributed.
For verification, we employed proprietary LLMs from OpenAI and enabled the opt-out setting to ensure that submitted data were not used for model training. For downstream applications involving personal or sensitive data, we recommend deploying open-source LLMs in a local environment to reduce the risk of unintended disclosure. The resources introduced in this work are compatible with locally hosted LLMs. The CEFR levels provided by our resources are model-based estimates and should not be used as the sole basis for high-stakes educational decisions, such as promotion, pass/fail determinations, or selective admissions.